OpenStudy (anonymous):

using subtraction formulas for sin and cos, verify cos (pi/2-theta)=sin theta and sin(pi/2-theta)=cos theta

OpenStudy (missinglink):

cos(A-B) = (cos(A)*cos(B)) + (sin(A)*sin(B)) cos(pi/2 - x) = cos(pi/2)*cos(x) + sin(pi/2)*sin(x) = 0*cos(x) + 1*sin(x) = sin(x) proving sin(pi/2-x) = cos(x) is similar to the method shown above but the formula for sin(A-B) = sin(A)cos(B) + cos(A)sin(B).

OpenStudy (anonymous):

oh lol so then i'd just do the same for sin(pi/1-theta)=cos theta.

OpenStudy (missinglink):

yep and make sure to use the right formula.

OpenStudy (anonymous):

sin(u-v)=sin(u)cos(v)-cos(u)sin(v)

OpenStudy (missinglink):

lol yes, I think I provided you the wrong formula for sin(A-B)

OpenStudy (anonymous):

its supposed to be subtract but i get it thank you:)

OpenStudy (missinglink):

You are welcome :)